Oceangirl:

A 5-pound bag of apples cost $3.50. Use that rate, to complete the table to identify equivalent ratios, and then answer the questions. Apples ? 5 *? 15 Cost 0.70 3.50 7 ?

swinn:

Let x represent the cost of 7 pounds of apples 5/3.50=7/x Cross multiplying to solve 5x = 7*3.50 x =